9.
Find the equation of the line whose x-intercept is 8 and whose y-intercept is -3, in standard form.

A. 8x - 3y = 24
B. 3x - 8y = 24
C. 3x + 8y = 24
D. -8x + 3y = 24

Answer by stanbon(75887)

Find the equation of the line whose x-intercept is 8 and whose y-intercept is -3, in standard form.
------
x/8 - y/3 = 1
----
Multply thru by 24 to get:
3x - 8y = 24
================
Cheers,
Stan H.
===================
B. 3x - 8y = 24